# ActSonic: Everyday Activity Recognition on Smart Glasses using Active Acoustic Sensing

### v1

@@ 5. User Study
In this section, we present a comprehensive overview of the user studies conducted to assess the performance of ActSonic. The objective of these studies is to evaluate the activity recognition pipeline under naturalistic conditions. To achieve this goal, we devised a diverse set of everyday activities to monitor throughout the study, recruited participants, and carried out both a semi-in-the-wild user study and a more extended fully in-the-wild study, both conducted in participants’ homes in an unconstrained environment.
@@ 5.1. Design of Activity Set
  • Bathroom (6 activities): rinse_mouth, brushing, flossing, brush_hair, flush_toilet, open_door
  • Kitchen and Dining Area (8 activities) : washing_hands, eating, drinking, pickup/putdown, pouring, chopping, wiping_surface, stirring
  • Bedroom and Living Area (13 activities): stationary, walking, sitting, coughing, yawning, talking, putting_on_outerwear, vacuum/cleaning_floor, throwing, stretching, using_phone/tab, squat, reading_book
@@ 5.2. Participants and Study Schedule
The ActSonic user studies received approval from the Institutional Review Board for Human Participant Research (IRB) at our organization. We enlisted 12 participants for a semi-in-the-wild study and 7 participants for a fully-in-the-wild study. Among the total 19 participants, with an average age of $(24.737\pm 3.445)$ years, ranging from 21 to 31 years, 5 identified as female, and 14 identified as male. Per IRB guidelines, each study lasted no longer than 2 hours (120 minutes), and participants received \$30 USD as compensation for their time. Post-study, we gathered basic demographic and physical information (e.g., height, weight, gender), along with general feedback on the ActSonic wearable device via an IRB-approved questionnaire.
@@ 5.4. Study Design
We conducted a 12-participant semi-in-the-wild study followed by an in-the-wild study with 7 participants. Both of these studies were conducted at participants’ homes in unconstrained settings. The design of the study protocols is discussed below.
@@ 5.4.1. Study - 01: Semi-in-the-wild User Study
5.4.2. Study - 02: In-the-wild User Study at Participants’ Home
We designed a longer-duration in-the-wild study with the goal of evaluating the ActSonic activity recognition system in the wild for extended hours. Furthermore, the activity set of ActSonic (listed in Subsec. 5.1) contains a total of 27 activities, including the null class, and it is unlikely to get samples of all those activities in the second segment (30 minutes) of the semi-in-the-wild study in a naturalistic setting. Since the IRB protocol allows a maximum duration of two hours for a single study, we designed this in-the-wild study with a duration of two hours. However, the protocol allows multiple studies with the same participant, and therefore the participants were given an option to take part in multiple consecutive studies. One out of the 7 participants opted for that choice and participated in two consecutive studies which were four hours long in total, and they were compensated twice. Hence, we accumulated 16 hours of in-the-wild data from this study.
We followed the same data collection procedure for the in-the-wild study as the second segment of the semi-in-the-wild study. One difference to be noted is that the experimenter left the participants’ homes after briefing them and setting up the data collection system. The participants returned the data collection system after two hours. In this in-the-wild study, the participants were not instructed with any specifics of the activities to be performed during the study; rather, they were allowed to continue their regular schedule at their homes. Note that we did not permit participants to leave their homes due to legal restrictions on video recording in public places in the country where the study was conducted.
@@ 6.1. Leave-one-participant-out Evaluation of Prompted Sessions
We conducted a leave-one-participant-out cross-validation evaluation on prompted sessions of the semi-in-the-wild study. In the initial segment of the semi-in-the-wild study (as described in Sec 5), involving 12 participants, each participant contributed $8$ sessions. We trained $12$ user-independent models, where, for instance, the model for participant P01 was trained solely on data from P02-P12 and tested on P01, and the same process was repeated for the other participants. The average macro F1-score for each participant ranged from $0.90$ to $0.95$, exhibiting a standard deviation of $0.035$. Across all participants, the average macro F1-score in the leave-one-participant-out evaluation was $0.934$. When examining individual activities, we observed high accuracy (macro F1-score) for all activities, ranging from $0.88$ to $0.97$ across participants in the semi-in-the-wild study with a prediction frequency of 1 Hz. Compared to prior work(Lane et al., 2015; Mollyn et al., 2022), which evaluated the activity recognition in similar manner (albeit different activities), ActSonic has significantly better performance with a wider range of activities.
However, we also admitted that in this part of the study, participants performed activities following audio stimuli with the presence of a researcher at their home, which may reduce the variance of how they perform activities in real-world settings.
@@ 6.2. Evaluation of User Independent Model on Unconstrained Sessions
Figure 7. In-the-wild evaluation of the performance of different activities.
@@
Figure 8. Activity-wise number of participants for the unconstrained sessions: with activity labels on the $x$-axis and number of participants on the $y$-axis.
Our evaluation of the unconstrained sessions from both the semi-in-the-wild and in-the-wild studies involves two stages. In the initial stage, we take each model trained on prompted sessions and conduct a leave-one-participant-out evaluation (as described in Subsec. 6.1). Subsequently, we fine-tune these models using unconstrained data from the semi-in-the-wild study for other participants before assessing the model on the original participant. For instance, the P01 ”prompted” model, trained on P02-P12 ”prompted” supervision, undergoes fine-tuning using ”unconstrained” supervision from P02-P12, ensuring the exclusion of labels from P01 during model training.
Figure 9. Normalized confusion matrix of leave-one-participant-out user evaluation in in-the-wild sessions
In the second phase of the unconstrained session evaluation, we measure the performance using data from the in-the-wild study. To assess the unconstrained sessions of users P13 to P19 from the in-the-wild study, we first train a fine-tuned model using prompted session data of users P01 to P12 as supervision. Subsequently, we evaluate the model performance using data from individual users of P13-P19 as the test set.
The average macro F1-score and standard deviation for all actions listed in Sec 5.1 are reported in Figure 6. Specifically, Figure 6(a) and 6(b) present the average activity recognition performance on semi-in-the-wild and in-the-wild participants respectively. Furthermore, Figure 7 displays the average macro F1-score and standard deviation for each of the in-the-wild actions across all participants. Our findings reveal an average F1-score of $0.866$ with a standard deviation of $0.052$ with a predication frequency of 1Hz. Although the system’s performance in an in-the-wild scenario is comparatively lower than in prompted sessions, it is apparent in Figure 9 that the activity recognition remains accurate, considering the user variability and class imbalance present in the in-the-wild sessions.
The class distribution, indicating the number of participants performing specific activities, is illustrated in Figure 8. The observed accuracy imbalance among different actions within the in-the-wild examples likely stems from the natural variability in individual interactions within an open-world setting, influenced by external context. While activities like ”rinsing mouth,” ”brushing,” ”walking,” and ”sitting/standing” seem less contextually dependent, actions such as ”pouring,” ”throwing,” and ”pickup/putdown” might display lower performance due to subtler movements and inherent variability arising from interactions with different objects.
The reported performance is a significant advancement in wearable-based activity recognition in the wild, as it is the first wearable-based (non-camera) activity recognition system that was evaluated in such an unconstrained environment, with a wide variety of activities in a user-independent manner, while lasting over 20 hours.
@@ 6.4. Latency and System Overhead of ActSonic on Mobile Platform
We evaluated the ActSonic system’s latency and overhead on the Google Pixel 7 mobile platform. Table 1 presents the inference time and various parameters of the ActSonic ResNet18 model evaluated on an in-the-wild dataset with the same protocol described in 6.2. Initially, we generated lightweight mobile models from both the original and the 8-bit integer quantized versions of the ResNet18 model using PyTorch Mobile. Subsequently, we conducted inference time benchmarks using a two-second sliding window on a Google Pixel 7 Android phone, performing inference on 1000 samples. The mean inference time is detailed in Table 1. The ActSonic performance on mobile devices was evaluated by transmitting the acoustic data to a Pixel 7 for inference. The BLE status was set to ”connected” since it continuously streams data to mobile devices.
Furthermore, we utilized the Android Profiler (Developers, [n. d.]) to evaluate mobile CPU usage and energy consumption. Table 1 indicates that while the quantized model exhibits lower accuracy compared to its non-quantized counterpart, it demonstrates lower system overhead. This assessment utilized a post-training quantization strategy; employing a quantization-aware training approach for the ResNet18 model might yield improved performance while preserving similar efficiency.
@@ 7.2. Performance Comparison of Different Deep Learning Encoders
Prompted Sessions In-the-wild Sessions
@@
Observing Table 2, we note that while self-supervision doesn’t exhibit significant performance improvement in the controlled sessions, it does demonstrate an impact in maintaining performance in variable in-the-wild scenarios. Furthermore, in comparison to the number of parameters of MobileNetV2, ResNet18 has a larger memory footprint. This observation is particularly valuable in the scenarios involving the performance-inference time tradeoff of the ActSonic system. Additionally, ConvLSTM exhibits worse performance compared to self-supervised ResNet18 despite having the explicit capacity to model temporal dependency.
@@ 7.3. Comparison of Performance with Prior Systems
Table 3. Comparison of performance of activity recognition systems with similar sensing modality. Note that the systems were not evaluated on the same dataset. Therefore, numerical differences may not provide a fair comparison. We also provide user study evaluation information in the table.
We compare ActSonic with other activity recognition systems in Table 3. Many previous methods primarily focus on action recognition using passive data, such as passive acoustics, passive ultrasonic/infrasonic sensing, egocentric camera data, or inertial measurements from smartwatch data. However, these methods may not capture all actions comprehensively. Furthermore, most previous studies concentrate on smaller-scale experiments with specialized sensors or large-scale video-only benchmarks, with few considering in-the-wild settings.
ActSonic’s utilization of an active acoustic sensing mechanism enables the capture of signals representing fine-grained movements that passive systems may miss. For instance, it can recognize actions occurring outside the frame of an egocentric camera or actions with minimal audio cues, which passive acoustic sensing may struggle to detect accurately. Additionally, passive acoustic sensing-based methods are vulnerable to changes in environmental parameters, hindering their ability to generalize signals across different environments and making deployment in the wild challenging.
As shown in Table 3, most passive sensing approaches can detect audio events such as microwave usage, blender operation, or alarm clock sounds. However, these events do not necessarily imply that the user of the wearable system is engaged in specific activities. In contrast, ActSonic can recognize a wide variety of fine-grained body motion-based activities compared to other systems, achieving over 90% accuracy in naturalistic settings. Conversely, existing systems, as observed in Table 3, often demonstrate similar or worse performance, even with a smaller activity set or in controlled settings.
@@ 8. Discussion
8.1. Saliency Analysis to Visualize Class Activation Feature Maps
@@
Figure 11. Selected GradCAM heatmaps overlaid on differential echo profiles. As we have a 4-channel input and GradCAM aggregates heat maps by channel, we overlay the same (smoothed) heatmap across all four channels. Redder values from the smooth interpolation correspond to higher significance towards class prediction, while bluer values indicate lower significance.
Our user study demonstrates competitive accuracy in a wearable-based activity recognition system utilizing active acoustic sensing via echo profiles. Additionally, we conduct a saliency analysis to identify significant sections of the echo profile input for specific classes. Employing Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019), we visualize the final convolutional layer of the ResNet18 encoder. Overlaying the feature heatmap on the 4 channels of acoustic flow or differential echo profiles confirms our model’s capability to capture class-related motion within these channels. Figure 11 illustrates Grad-CAM for randomly picked sliding windows of all activities included in the ActSonic set.
In Figure 11, distinct activities exhibit activation in various regions of the echo profile sliding window. This observation validates that ActSonic’s self-supervised model learns to focus on different explainable regions within the input sliding window to infer everyday activities. For instance, in the activity of brushing teeth, repetitive movement near the face region is evident, with the ResNet18 encoder displaying higher gradient values (indicative of heightened attention) in that area to infer the activity. Similarly, the activity of coughing illustrates hand movements in front of the face and the resulting motion pattern induced in the face due to coughing. Additionally, the activity of opening a door demonstrates hand movement to unlock the door, followed by movement to enter the room.

### v2

@@ 5. User Study
In this section, we present a comprehensive overview of the user studies conducted to assess the performance of ActSonic. The objective of these studies is to evaluate the activity recognition pipeline under naturalistic conditions. To achieve this goal, we devised a diverse set of everyday activities to monitor throughout the study, recruited participants, and carried out both a semi-in-the-wild user study and a more extended unconstrained study, both conducted in participants’ homes in a naturalistic environment.
@@ 5.1. Design of Activity Set
  • Bathroom (6 activities): Rinse Mouth, Brushing Teeth, Flossing, Brushing Hair, Flush Toilet, Opening Door
  • Kitchen and Dining Area (8 activities) : Washing Hands, Eating, Drinking, Pickup / Putdown, Pouring, Chopping, Wiping Surface, Stirring
  • Bedroom and Living Area (13 activities): Stationary, Walking, Sitting, Coughing, Yawning, Talking, Putting on Outerwear, Vacuum, Throwing, Stretching, Using Phone/Tab, Squat, Reading Book
@@ 5.2. Participants and Study Schedule
The ActSonic user studies received approval from the Institutional Review Board for Human Participant Research (IRB) at our organization. We enlisted 12 participants for a semi-in-the-wild study and 7 different participants for an unconstrained user study. Among the total 19 participants, with an average age of $(24.737\pm 3.445)$ years, ranging from 21 to 31 years, 5 identified as female, and 14 identified as male. Per IRB guidelines, each study lasted no longer than 2 hours (120 minutes), and participants received \$30 USD as compensation for their time. Post-study, we gathered basic demographic and physical information (e.g., height, weight, gender), along with general feedback on the ActSonic wearable device via an IRB-approved questionnaire.
@@ 5.4. Study Design
We conducted a 12-participant semi-in-the-wild study followed by an unconstrained study with 7 participants. Both of these studies were conducted at participants’ homes in unconstrained settings. The design of the study protocols is discussed below.
5.4.1. User Study Design Considerations
We conducted two separate studies at participants’ homes in a naturalistic setting. Given that the activity set of ActSonic encompasses a wide range of activities, obtaining samples of each activity in an uncontrolled setting, where participants are allowed to carry out their daily routines, proved challenging. Therefore, we conducted a semi-in-the-wild study where users were prompted with specific instructions to perform activities, thus covering the tracking set of ActSonic. Subsequently, we conducted a second unconstrained study at participants’ homes, where they did not receive any particular instructions to perform activities but instead continued with their daily routines.
We utilized a chest-mounted camera to gather reference videos for ground truth annotation. In contrast to other activity recognition systems evaluated in real-world settings (Shin et al., 2022), we opted not to rely on self-reporting-based annotation, despite its facilitation of longer duration in-the-wild data collection without significant labeling effort. This decision stems from the impracticality of accurately self-reporting all 27 types of activities throughout a day. Additionally, we aimed to evaluate ActSonic based on the inference generated at each second, necessitating fine-grained ground truth annotation from the uncontrolled study, which is only achievable with a body-mounted camera.
@@ 5.4.2. Study - 01: Semi-in-the-wild User Study
5.4.3. Study - 02: Unconstrained Study at Participants’ Home
We designed a longer-duration unconstrained study with the goal of evaluating the ActSonic activity recognition system in the wild for extended hours. Furthermore, the activity set of ActSonic (listed in Subsec. 5.1) contains a total of 27 activities, including the null class, and it is unlikely to get samples of all those activities in the second segment (30 minutes) of the semi-in-the-wild study in a naturalistic setting. Since the IRB protocol allows a maximum duration of two hours for a single study, we designed this in-the-wild study with a duration of two hours. However, the protocol allows multiple studies with the same participant, and therefore the participants were given an option to take part in multiple consecutive studies. One out of the 7 participants opted for that choice and participated in two consecutive studies which were four hours long in total, and they were compensated twice. Hence, we accumulated 16 hours of in-the-wild data from this study.
We followed the same data collection procedure for the unconstrained study as the second segment of the semi-in-the-wild study. One difference to be noted is that the experimenter left the participants’ homes after briefing them and setting up the data collection system. The participants returned the data collection system after two hours. In this unconstrained study, the participants were not instructed with any specifics of the activities to be performed during the study; rather, they were allowed to continue their regular schedule at their homes. Note that we did not permit participants to leave their homes due to legal restrictions on video recording in public places in the country where the study was conducted.
@@ 6.1. Leave-one-participant-out Evaluation of Prompted Sessions
We conducted a leave-one-participant-out cross-validation evaluation on prompted sessions of the semi-in-the-wild study. In the initial segment of the semi-in-the-wild study (as described in Sec 5), involving 12 participants, each participant contributed $8$ sessions. We trained $12$ user-independent models, where, for instance, the model for participant P01 was trained solely on data from P02-P12 and tested on P01, and the same process was repeated for the other participants. The average macro F1-score for each participant ranged from $0.90$ to $0.95$, exhibiting a standard deviation of $0.035$. Across all participants, the average macro F1-score in the leave-one-participant-out evaluation was $0.934$. When examining individual activities, we observed high accuracy (macro F1-score) for all activities, ranging from $0.88$ to $0.97$ across participants in the semi-in-the-wild study with a prediction frequency of 1 Hz. Compared to prior work(Lane et al., 2015; Mollyn et al., 2022), which evaluated activity recognition in a similar manner (albeit with different activities), ActSonic has significantly better performance with a wider range of activities. However, we also acknowledge that in this part of the study, participants performed activities following audio stimuli with the presence of a researcher at their home, which may reduce the variance in how they perform activities in real-world settings.
@@ 6.2. Evaluation of User Independent Model on Unconstrained Sessions
Figure 7. Evaluation of the performance of different activities in unconstrained sessions
@@
Figure 8. Number of participants per activity during unconstrained sessions: activity labels on the $x$-axis and number of participants on the $y$-axis
Our evaluation of the unconstrained sessions from both the semi-in-the-wild and the second studies involves two stages. In the initial stage, we take each model trained on prompted sessions and conduct a leave-one-participant-out evaluation (as described in Subsec. 6.1). Subsequently, we fine-tune these models using unconstrained data from the semi-in-the-wild study for other participants before assessing the model on the original participant. For instance, the P01 ”prompted” model, trained on P02-P12 ”prompted” supervision, undergoes fine-tuning using ”unconstrained” supervision from P02-P12, ensuring the exclusion of labels from P01 during model training.
Figure 9. Normalized confusion matrix of leave-one-participant-out user evaluation in unconstrained sessions
In the second phase of the unconstrained session evaluation, we measure the performance using data from the second study. To assess the unconstrained sessions of users P13 to P19 from the unconstrained study, we first train a fine-tuned model using prompted session data of users P01 to P12 as supervision. Subsequently, we evaluate the model performance using data from individual users of P13-P19 as the test set.
The average macro F1-score and standard deviation for all actions listed in Sec 5.1 are reported in Figure 6. Specifically, Figure 6(a) and 6(b) present the average activity recognition performance on semi-in-the-wild and unconstrained study participants respectively. Furthermore, Figure 7 displays the average macro F1-score and standard deviation for each of the unconstrained actions across all participants. Our findings reveal an average F1-score of $0.866$ with a standard deviation of $0.052$ with a predication frequency of 1 Hz. Although the system’s performance in an unconstrained scenario is comparatively lower than in prompted sessions, it is apparent in Figure 9 that the activity recognition remains accurate, considering the user variability and class imbalance present in the unconstrained sessions.
The class distribution, indicating the number of participants performing specific activities, is illustrated in Figure 8. The observed accuracy imbalance among different actions within the unconstrained examples likely stems from the natural variability in individual interactions within an open-world setting, influenced by external context. While activities like ”rinsing mouth,” ”brushing,” ”walking,” and ”sitting/standing” seem less contextually dependent, actions such as ”pouring,” ”throwing,” and ”pickup/putdown” might display lower performance due to subtler movements and inherent variability arising from interactions with different objects.
@@ 6.4. Latency and System Overhead of ActSonic on Mobile Platform
We evaluated the latency and overhead of the ActSonic system on the Google Pixel 7 mobile platform. Table 1 presents the inference time and various parameters of the ActSonic ResNet18 model evaluated on an unconstrained dataset with the same protocol described in 6.2. Initially, we generated lightweight mobile models from both the original and the 8-bit integer quantized versions of the ResNet18 model using PyTorch Mobile. Subsequently, we conducted inference time benchmarks using a two-second sliding window on a Google Pixel 7 Android phone, performing inference on 1000 samples. The mean inference time is detailed in Table 1. The ActSonic performance on mobile devices was evaluated by transmitting the acoustic data to a Pixel 7 for inference. The BLE status was set to ”connected” since it continuously streams data to mobile devices.
Furthermore, we utilized the Android Profiler (Developers, [n. d.]) to evaluate mobile CPU usage and energy consumption. Table 1 indicates that while the quantized model exhibits lower accuracy compared to its non-quantized counterpart, it demonstrates lower system overhead. This assessment utilized a post-training quantization strategy; employing a quantization-aware training approach for the ResNet18 model might yield improved performance while preserving similar efficiency. The energy consumption was indicated as ”light” for both the quantized and non-quantized models by the Android Profiler (Developers, [n. d.]).
@@ 7.2. Performance Comparison of Different Deep Learning Encoders
Prompted Sessions Unconstrained Sessions
@@
Observing Table 2, we note that while self-supervision doesn’t exhibit significant performance improvement in the controlled sessions, it does demonstrate an impact in maintaining performance in variable unconstrained scenarios. Furthermore, in comparison to the number of parameters of MobileNetV2, ResNet18 has a larger memory footprint. This observation is particularly valuable in the scenarios involving the performance-inference time tradeoff of the ActSonic system. Additionally, ConvLSTM exhibits worse performance compared to self-supervised ResNet18 despite having the explicit capacity to model temporal dependency.
@@ 7.3. Comparison of Performance with Prior Systems
Table 3. Comparison of performance of activity recognition systems with similar sensing modality. Note that the systems were not evaluated on the same dataset. Therefore, numerical differences may not provide a fair comparison. User study evaluation information is also provided in the table.
We compare ActSonic with other activity recognition systems in Table 3. ActSonic, an active acoustic sensing-based activity recognition system, is compared with systems utilizing audible acoustic data, passively sensed inaudible (ultrasonic and infrasonic) acoustic data, motion data (captured through IMU), and egocentric camera data.
Activity recognition systems rely on passively sensed audible signals (such as BodyScope (Yatani and Truong, 2012), Ubicoustics (Laput et al., 2018)) or inaudible acoustics (PrivacyMic (Iravantchi et al., 2021)). These systems operate on the premise that different actions generate distinct acoustic signatures in the environment. Consequently, their activity sets include actions like chopping, baby crying, or using a mixer. However, many daily activities, such as wearing clothes, reading, or flossing, do not produce discernible sounds suitable for activity inference. Moreover, these activities may not necessarily involve the user wearing the device, especially with wearables. Additionally, previous studies have mainly focused on smaller-scale experiments with specialized sensors or large-scale video-only benchmarks (e.g., DiffAct (Liu et al., 2023)), with few considering real-world settings.
On the other hand, ActSonic can track fine-grained actions such as flossing or wiping surfaces since its sensing mechanism relies on tracking the movement of different body parts simultaneously. It is evident from Table 3 that ActSonic achieves higher recognition performance (Macro F1-score over $90\%$) in naturalistic environments, with an extensive set of fine-grained everyday activities. ActSonic’s utilization of an active acoustic sensing mechanism enables the capture of signals representing fine-grained movements that passive sensing-based systems may miss. For instance, it can recognize actions occurring outside the frame of an egocentric camera or actions with minimal audio cues, which passive acoustic sensing may struggle to detect accurately. Additionally, passive acoustic sensing-based methods are vulnerable to changes in environmental parameters, hindering their ability to generalize signals across different environments and making deployment in the wild challenging.
@@ 8. Discussion
8.1. Visual Explanation of Learned Parameters of ActSonic Model through Saliency Analysis
@@
Figure 11. Selected Grad-CAM (Selvaraju et al., 2019) heatmaps overlaid on differential echo profiles. As we have a 4-channel input and Grad-CAM aggregates heat maps by channel, we overlay the same (smoothed) heatmap across all four channels. Redder values from the smooth interpolation correspond to higher significance towards class prediction, while bluer values indicate lower significance. We display a SMPL (Loper et al., 2023) mesh of the body region covered by the sensing range of ActSonic’s sliding window input in the rightmost column.
Our user study demonstrated competitive accuracy in recognizing 27 everyday activities by learning patterns from reflected inaudible acoustic waves around the body. The reflected acoustic signals are formed with complex multipath echoes from both the body and the surrounding environments. Additionally, the environment can occasionally generate acoustic signals with frequency components that overlap with our emitted frequency range. Therefore, received acoustic signals contain rich information about the environment in addition to the body poses.
A common critique of many data-driven sensing systems is that machine learning algorithms operate as a black box, making it difficult to understand what and how they learn. Given the complex reflections forming the received acoustic signals, it is critically important to understand what the deep learning algorithm actually learns from these signals.
To address this, we conducted saliency analysis using Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019) to visualize the ResNet18 encoder’s final convolutional layer. Overlaying feature heatmaps on acoustic flow or differential echo profile channels confirm the model’s ability to capture the movements of body parts and areas related to each activity. Figure 11 demonstrates Grad-CAM (Selvaraju et al., 2019) for all activities in the dataset.
As shown in Figure 11, various activities activate distinct regions within the echo profile sliding window in the neural networks. The $y$-axis represents the distance from reflected echoes to the sensor, while the $x$-axis denotes the timeline. To help comprehension, we aligned the $y$-axis with a human body model. Colors in the figure indicate body movement intensity, with red indicating high intensity and dark blue indicating lower intensity.
The visualization clearly shows that the deep learning algorithms allocate significantly higher attention to areas with heavy body movements (e.g., face, chest, arms) during different activities. For instance, in the activity of brushing teeth, repetitive movement near the face region is evident, with the ResNet18 encoder displaying higher gradient values (indicative of heightened attention) in that area around the face ($5$ cm-$17.15$ cm) to infer the activity.
While some activities share similar arm movement patterns, fine details in reflected acoustic waves around the face differentiate them. For instance, both drinking and eating involves moving the arm toward the mouth. However, after feeding during eating, the hand usually moves downward, leaving only mouth movements, visible as subtle highlights under $10$ cm. Conversely, during drinking, the hand and container (e.g., cup) remain around the mouth, evident as a large area of heated sections ($15$-$30$ cm). Additionally, the activity of opening a door demonstrates hand movement to unlock the door, followed by movement to enter the room.
This visualization confirms that our self-supervised model focuses on different explainable regions of the echo profile within the input sliding window to infer everyday activities. It underscores the importance of capturing detailed movements in both the upper body and face areas to distinguish fine-grained activities such as eating versus drinking, which may not be distinguished in isolation.
8.2. Deploying ActSonic on Other Wearable Form Factors
ActSonic employs an active acoustic sensing system integrated into eyeglasses to emit inaudible acoustic signals around the body and capture the reflected waves. Due to the low-power and compact nature of the microphones and speakers, this technology can be easily adapted to other wearable form factors for tracking human activities using a similar sensing principle. Here, we evaluate the feasibility of implementing ActSonic on different form factors. Two critical considerations for applying ActSonic to wearable form factors are: 1) the ability to continuously emit acoustic waves around both sides of the upper body and head, and 2) the stability of the wearable form factor during various activities.
Based on these criteria, we anticipate that ActSonic can be seamlessly deployed with minimal modifications on wearable form factors worn around the head (pending training data collection for each form factor), such as necklaces, earbuds, or the recently introduced Humane AI Pin (Humane, [n. d.]) worn on the chest. These form factors provide similar stability to eyeglasses while continuously covering both the upper body and head to track detailed facial and upper body movements.
Wearable form factors attached to limbs, such as wristbands, armbands, or rings, pose greater challenges for applying ActSonic for activity recognition with a single device. Limbs exhibit a wide range of movements, complicating signal interpretation as the source moves concurrently with other body parts. Additionally, signals emitted from one side of the limb may be obstructed by the body, hindering sensing on both sides. Moreover, capturing facial movements from limb-mounted devices is challenging due to the relatively long distance and potential blockages. These issues may be mitigated with updated algorithms and hardware designs, which we plan to explore in future research.